40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Oct 30 01:56:38 2013, tara, DailyProgress, optic, table work oldPBS.jpgnewPBS.jpg
    Reply  Thu Oct 31 00:02:17 2013, tara, DailyProgress, Electronics Equipment, TTFSS offset.pngD040105C.png
       Reply  Thu Oct 31 18:10:07 2013, Evan, DailyProgress, Electronics Equipment, TTFSS 
       Reply  Sun Nov 3 19:37:32 2013, rana, DailyProgress, Electronics Equipment, TTFSS 
          Reply  Tue Nov 5 01:11:15 2013, tara, DailyProgress, Electronics Equipment, TTFSS 
             Reply  Wed Nov 6 01:14:58 2013, tara, DailyProgress, Electronics Equipment, TTFSS beat_2013_11_05.pngbeat_2013_11_05.fig
                Reply  Thu Nov 7 05:08:13 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_2013_11_07.pngbeat_2013_11_07.fig
                   Reply  Wed Nov 20 18:19:01 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_compare.pngbeat_compare.fig
Message ID: 1383     Entry time: Wed Nov 6 01:14:58 2013     In reply to: 1381     Reply to this: 1384
Author: tara 
Type: DailyProgress 
Category: Electronics Equipment 
Subject: TTFSS 

We made a mistake by choosing the input power to the cavities to be 0.25 mW, so today I turned them back to 1mW and measure the beat.

beat_2013_11_05.png

Setup:

  • input power: 1mW
  • TTFSS gain (C/F): RCAV (760/980), ACAV (630/650)

Note about the measurement:

  • The noise at error point from ACAV is pretty high (~up to 100 nV/rtHz around 6 kHz). Better characterization will be done later to see if the suppression is enough or not. I made sure that this measurement is good up to ~2kHz, (this was done by changing the gain level a bit and beat level did not change).
  • Table was floated, but the air springs was not activated. I hope we will get better signal around 20-100 Hz once the air spring is re-installed.
  •  Intensity noise->photothermal around a few hundred mHz, cause the PLL to drift away from the input range. This effect becomes worse when the input power is increased.
  • RIN from ACAV is about a factor of 10 higher than that of RCAV.
  • I measured the beat with ISS on/off on ACAV, nothing were significantly different. So maybe it is not a problem for now.

To do next:

  • I'll compare this measurement with the one from 8" cavity to see if the results agree or not. They are different mirrors, but from the same coating run. I suspect that the loss in these mirrors are higher than what we estimate (Ta2O5=2.5e-4, SiO2 1e-4).
  • Think about the error in the measurement (calibration, spotsize etc).
Attachment 2: beat_2013_11_05.fig  187 kB
ELOG V3.1.3-